# Characterizing higher Auslander(-Gorenstein) Algebras

### v1

@@ 4. Cotorsion Pairs
As mentioned in the introduction, this section is devoted to study the relation between higher Auslander-Gorenstein and cotorsion pairs and we present some results on higher Auslander-Gorenstein algebras from the viewpoint of cotorsion pairs and relate them to the notion of torsion-cotorsion triples introduced by Bauer, Botnan, Oppermann and Steen [BBOS].
@@
is a hereditary cotorsion pair in ${\rm{{mod}}}(\Gamma)$, where $\rm{dom}^{\geq i}(\Gamma)$ denotes the full subcategory of ${\rm{{mod}}}(\Gamma)$ consisting of all finitely generated $\Gamma$-modules having the dominant dimension at least $i$ and $\mathrm{inj}^{\leq i}(\Gamma)$ denotes the full subcategory of ${\rm{{mod}}}(\Gamma)$ consisting of all finitely generated $\Gamma$-modules with injective dimension at most $i$, see [Ma92, Lemma 3, Proposition 5, and the corollary before it].
A result by Marczinzik shows that ${\mathrm{Gproj}}^{\leq n+1-i}(\Gamma)=\rm{dom}^{\geq i}(\Gamma)$ (see, [Ma18, Theorem 2.1]) and so, by [ET01, Theorem 10] or Salce’s lemma and [Ma92, Proposition 7], the pair
@@
Let $n$ be a positive integer, $10\leq k\leq n+$ 1, $\Gamma$ be an $n$-minimal Auslander-Gorenstein algebra; and $Q$ be its maximal projective-injective summand. Then $T_{k}:=\Omega^{-k}(\Gamma)\oplus Q$ is a $k$-tilting module and ${\rm{add}}(T_{k})={\mathrm{Gproj}}^{\leq k}(\Gamma)\cap\mathrm{inj}^{\leq n+1-k}(\Gamma)$. Also for all $0\leq j\leq n+1-k$,
@@
be the minimal injective resolution of $\Gamma$. Since $I^{0},I^{1},\cdots,I^{n}\in{\rm{add}}(\mathcal{Q})$, clearly, $\Gamma$ has an ${\rm{add}}(T_{k})$-copresentation of length $k$ and for every $10\leq k\leq n+$ 1, $\Omega^{-k}(\Gamma)\in{\mathrm{Gproj}}^{\leq k}(\Gamma)\cap\mathrm{inj}^{\leq n+1-k}(\Gamma)$. Hence, for every $11\leq l\leq n+$ 1,
@@
Therefore, $T_{k}$ is a $k$-tilting module; see [HU96, Page 448] or [AC01, Section 2] for the definition.
@@
To complete the proof, we must show that for every $10\leq k\leq n+$ 1, ${\rm{add}}(T_{k})={\mathrm{Gproj}}^{\leq k}(\Gamma)\cap\mathrm{inj}^{\leq n+1-k}(\Gamma)$. Clearly, ${\rm{add}}(T_{k})\subseteq{\mathrm{Gproj}}^{\leq k}(\Gamma)\cap\mathrm{inj}^{\leq n+1-k}(\Gamma)$ and so we only need to prove the converse.
Let $X\in{\mathrm{Gproj}}^{\leq k}(\Gamma)\cap\mathrm{inj}^{\leq n+1-k}(\Gamma)$. By [Ho04b, Theorem 2.2], ${\mathrm{Gproj}}^{\leq k}(\Gamma)\cap\mathrm{inj}^{\leq n+1-k}(\Gamma)=\mathrm{proj}^{\leq k}(\Gamma)\cap\mathrm{inj}^{\leq n+1-k}(\Gamma)$ and so ${\rm{pd}}(_{\Gamma}X)\leq k$. Hence, by [AC01, Lemma 2.2 $(a)$], ${}^{\perp}(T_{k}^{\perp})\subseteq\mathrm{proj}^{\leq k}(\Gamma)\subseteq{\mathrm{Gproj}}^{\leq k}(\Gamma)$ and so
@@
On the other hand, by [AC01, Lemmas 2.3 $(a)$ and 2.4 $(a)$], for every $Y\in T_{k}^{\perp}$, there is a short exact sequence
@@
where $T^{\prime}_{0},\cdots,T^{\prime}_{k-1}\in{\rm{add}}(T_{k})$. Clearly $K_{k-1}\in X^{\perp_{[k+1,\infty)}}$ and $T^{\prime}_{0},\cdots,T^{\prime}_{k-1}\in{\rm{add}}(T_{k})\subseteq X^{\perp}$. Hence, a dimension shifting argument shows that $Y\in X^{\perp}$. Therefore, $X\in{{}^{\perp}(T_{k}^{\perp})}$. But by [AC01, Lemmas 2.3 $(a)$ and 2.4 $(a)$] there is a short exact sequence
@@
Among the above mentioned cotorsion pairs, one of them has a very nice characteristic and its cotorsion class is a torsion-free class of a hereditary torsion pair. This fact relates higher Auslander-Gorenstein Algebras to the notion of torsion-cotorsion triples introduced by Bauer, Botnan, Oppermann and Steen [BBOS] and gives us a nice description of them.
@@
A torsion cotorsion triple in an abelian category $\mathcal{A}$ is a triple of subcategories $(\mathcal{T},\mathcal{F},\mathcal{D})$ such that the pair $(\mathcal{T},\mathcal{F})$ is a torsion pair and the pair $(\mathcal{F},\mathcal{D})$ is a cotorsion pair, see [BBOS, Page 29 before Theorem 2.33].
Theorem 4.3 (Dual of [BBOS, Theorem 2.35]).
@@
This theorem is the dual version of [BBOS, Theorem 2.35] and for the convenience of the reader we give the proof.
By [ARS, Proposition IV.1.9], we know that the inverse of Auslander-Reiten translation $\tau^{-}$ defines an equivalence
[…]
@@
where the middle equivalence follows from Auslander-Reiten formula because the injective dimension of every element of $\mathcal{D}$ by Bauer-Botnan-Oppermann-Steen Correspondence is at most one, see [SY11, Corollary III.6.4] and [BBOS, Theorem 2.34]. ∎
@@
Since $\Gamma$ is an $n$-minimal Auslander-Gorenstein algebra, the triple $(^{\perp_{0}}\Gamma,{\mathrm{Gproj}}^{\leq n}(\Gamma),\mathrm{inj}^{\leq 1}(\Gamma))$ is a torsion cotorsion triple, see Corollary 2.7 and the arguments at the beginning of this section. Hence, by [BBOS, Theorem 2.33], we have the first equivalence. That is
@@ 5. Higher Auslander Algebras
$\Gamma$ is also called a higher Auslander algebra, if there is a positive integer $n$ such that $\Gamma$ is an $n$-Auslander algebra. Iyama introduced these algebras as a generalization of classical Auslander algebras [Iy11, Page 3].
@@
Let $\Gamma$ be an Artin algebra and $n$ be a positive integer. Then we have the following statements:
  • $(a)$
    If $\mathrm{proj}(\Gamma)$ is an abelian category, then
    $\mathrm{proj}^{\leq 1}(\Gamma)\cap\mathrm{inj}(\Gamma)\subseteq\mathrm{proj}(\Gamma)\ \mathrm{and}\ {{}^{\perp_{0}}\Gamma}\subseteq{{}^{\perp_{1}}\Gamma};$
  • $(b)$
    If ${\mathrm{Proj}}(\Gamma)$ is an abelian category, then
    ${\mathrm{Proj}}^{\leq 1}(\Gamma)\cap{\mathrm{Inj}}(\Gamma)\subseteq{\mathrm{Proj}}(\Gamma)\ \mathrm{and}\ {{}^{\underline{\perp}_{0}}\Gamma}\subseteq{{}^{\underline{\perp}_{1}}\Gamma}.$
Hence, Theorem 5.1 is precisely a generalization of Auslander-Tachikawa theorem (Theorem 1) and also provide a new proof for it.
Let us prove $(a)$, that of $(b)$ being similar. In fact, assume that $\mathrm{proj}(\Gamma)$ is an abelian category. By Step 1 of the proof of Proposition 2.4, we see that $\mathrm{proj}(\Gamma)$ has the same kernels as ${\rm{{mod}}}(\Gamma)$, so $\mathrm{gldim}(\Gamma)\leq 2$.
Let $X\in\mathrm{proj}^{\leq 1}(\Gamma)\cap\mathrm{inj}(\Gamma)$. Since $\mathrm{proj}(\Gamma)$ is an abelian category, by the proof of Proposition 3.1, there exists a projective $\Gamma$-module $Y$ such that $X$ is a submodule of $Y$. But $X$ is an injective $\Gamma$-module and so it is a direct summand of $Y$. Therefore, $X$ is a projective $\Gamma$-module. This shows that $\mathrm{proj}^{\leq 1}(\Gamma)\cap\mathrm{inj}(\Gamma)\subseteq\mathrm{proj}(\Gamma)$.
Let now $X\in{{}^{{\perp}_{0}}}\Gamma$. Then, there is an exact sequence
$0\longrightarrow K\stackrel{{\scriptstyle\kappa}}{{\longrightarrow}}P_{1}\stackrel{{\scriptstyle\beta}}{{\longrightarrow}}P_{0}\longrightarrow X\longrightarrow 0$
where $P_{0},P_{1}\in\mathrm{proj}(\Gamma)$. Since $X\in{{}^{{\perp}_{0}}}\Gamma$, ${\rm{Hom}}_{\Gamma}(X,P)=0$ for every finitely generated projective $\Gamma$-module $P$, $\beta$ is a surjective map in $\mathrm{proj}(\Gamma)$. Since $\mathrm{gldim}(\Gamma)\leq 2$, $K$ is projective. Thus the sequence $0\longrightarrow K\stackrel{{\scriptstyle\kappa}}{{\longrightarrow}}P_{1}\stackrel{{\scriptstyle\beta}}{{\longrightarrow}}P_{0}\longrightarrow 0$ is exact in $\mathrm{proj}(\Gamma)$ and so by the exact sequence of abelian groups
$0\longrightarrow{\rm{Hom}}_{\Gamma}(P_{0},\Gamma)\longrightarrow{\rm{Hom}}_{\Gamma}(P_{1},\Gamma)\longrightarrow{\rm{Hom}}_{\Gamma}(K,\Gamma)$
one can easily see that ${\rm{Ext}}^{1}_{\Gamma}(X,\Gamma)=0$. Thus the condition ${{}^{{\perp}_{0}}}\Gamma\subseteq{{}^{{\perp}_{1}}}\Gamma$ holds.
Remark 5.3.
For $n\geq 2$, the conditions $\mathrm{proj}^{\leq n}(\Gamma)\cap\mathrm{inj}(\Gamma)\subseteq\mathrm{proj}(\Gamma)$ and ${{}^{\perp_{0}}\Gamma}\subseteq{{}^{\perp_{n}}\Gamma}$ in Theorem 5.1(b) are necessary. In fact, take an Artin algebra of global dimension $n-1$ which is not an $n$-Auslander algebra. Then $\mathrm{proj}^{\leq n-1}(\Gamma)={\rm{{mod}}}(\Gamma)$ is an abelian category, but by Theorem 5.1, the conditions that $\mathrm{proj}^{\leq n}(\Gamma)\cap\mathrm{inj}(\Gamma)\subseteq\mathrm{proj}(\Gamma)$ or ${{}^{\perp_{0}}\Gamma}\subseteq{{}^{\perp_{n}}\Gamma}$ don’t hold.
@@
[…]
@@
The following example gives a better perspective about some results of the paper. In the following, for every vertex $i$ of the given quiver $Q$, the corresponding simple, projective, and injective object will be deonted by $S(i)$, $P(i)$, and $I(i)$, respectively. For more details about quiver representations, see for example [SY11].
@@
    By [SY11, Theorem I.10.5], the Nakayama algebra $\Gamma$ is of finite representation type with $32n+$ 3 indecomposable $\Gamma$-modules. The following table gives us more information about their homological dimensions.
@@
    $0\longrightarrow P(1)\longrightarrow P(n+1)\longrightarrow S(n+1)\longrightarrow 0$
@@
Let $R$ be a commutative artinian ring and $n$ be a positive integer. In 2017, Iyama and Jasso extended higher Auslander correspondence from Artin $R$-algebras of finite representation type to dualizing $R$-varieties. By their result, i.e. higher Auslander correspondence for dualizing $R$-varieties, we know that an Artin $R$-algebra $\Gamma$ is an $n$-Auslander algebra if and only if the category of all finitely generated projective $\Gamma$-modules is an $n$-abelian category [IJ17, Theorem 1.2]. But by some results due to Beligiannis [Be15], even for $n$-minimal Auslander-Gorenstein algebras, one can see that the category of all finitely generated Gorenstein projective $\Gamma$-modules is not an $n$-abelian category in general. Hence, based on our results, there is a difference between the abelianness of ${\mathrm{Gproj}}^{\leq n-1}(\Gamma)$ and the $n$-abelianness of ${\mathrm{Gproj}}(\Gamma)$ in general.
Acknowledgements: The authors were supported by the National Natural Science Foundation of China (No. 12071137), by Key Laboratory of MEA (Ministry of Education), by Shanghai Key Laboratory of PMMP (No. 22DZ2229014), and by Fundamental Research Funds for the Central Universities.

### v2

@@ 5. Cotorsion Pairs
As mentioned in the introduction, this section is devoted to study the relation between higher Auslander-Gorenstein and cotorsion pairs and we present some results on higher Auslander-Gorenstein algebras from the viewpoint of cotorsion pairs and relate them to the notion of torsion-cotorsion triples introduced by Bauer, Botnan, Oppermann and Steen [11].
@@
is a hereditary cotorsion pair in ${\rm{{mod}}}(\Gamma)$, where ${\rm{dom}}^{\geq i}(\Gamma)$ denotes the full subcategory of ${\rm{{mod}}}(\Gamma)$ consisting of all finitely generated $\Gamma$-modules having the dominant dimension at least $i$ and $\mathrm{inj}^{\leq i}(\Gamma)$ denotes the full subcategory of ${\rm{{mod}}}(\Gamma)$ consisting of all finitely generated $\Gamma$-modules with injective dimension at most $i$, see [36, Lemma 3, Proposition 5, and the corollary before it].
A result by Marczinzik shows that ${\mathrm{Gproj}}^{\leq n+1-i}(\Gamma)={\rm{dom}}^{\geq i}(\Gamma)$ (see, [35, Theorem 2.1]) and so, by [17, Theorem 10] or Salce’s lemma and [36, Proposition 7], the pair
@@
Let $n$ be a positive integer, $0\leq k\leq n+1$, $\Gamma$ be an $n$-minimal Auslander-Gorenstein algebra; and $Q$ be its maximal projective-injective summand. Then $T_{k}:=\Omega^{-k}(\Gamma)\oplus Q$ is a $k$-tilting module and ${\rm{add}}(T_{k})={\mathrm{Gproj}}^{\leq k}(\Gamma)\cap\mathrm{inj}^{\leq n+1-k}(\Gamma)$. Also for all $0\leq j\leq n+1-k$,
@@
be the minimal injective resolution of $\Gamma$. Since $I^{0},I^{1},\cdots,I^{n}\in{\rm{add}}(\mathcal{Q})$, clearly, $\Gamma$ has an ${\rm{add}}(T_{k})$-copresentation of length $k$ and for every $0\leq k\leq n+1$, $\Omega^{-k}(\Gamma)\in{\mathrm{Gproj}}^{\leq k}(\Gamma)\cap\mathrm{inj}^{\leq n+1-k}(\Gamma)$. Hence, for every $1\leq l\leq n+1$,
@@
Therefore, $T_{k}$ is a $k$-tilting module; see [19, Page 448] or [2, Section 2] for the definition.
@@
To complete the proof, we must show that for every $0\leq k\leq n+1$, ${\rm{add}}(T_{k})={\mathrm{Gproj}}^{\leq k}(\Gamma)\cap\mathrm{inj}^{\leq n+1-k}(\Gamma)$. Clearly, ${\rm{add}}(T_{k})\subseteq{\mathrm{Gproj}}^{\leq k}(\Gamma)\cap\mathrm{inj}^{\leq n+1-k}(\Gamma)$ and so we only need to prove the converse.
Let $X\in{\mathrm{Gproj}}^{\leq k}(\Gamma)\cap\mathrm{inj}^{\leq n+1-k}(\Gamma)$. By [21, Theorem 2.2], ${\mathrm{Gproj}}^{\leq k}(\Gamma)\cap\mathrm{inj}^{\leq n+1-k}(\Gamma)=\mathrm{proj}^{\leq k}(\Gamma)\cap\mathrm{inj}^{\leq n+1-k}(\Gamma)$ and so ${\rm{pd}}(_{\Gamma}X)\leq k$. Hence, by [2, Lemma 2.2 $(a)$], $^{\perp}(T_{k}^{\perp})\subseteq\mathrm{proj}^{\leq k}(\Gamma)\subseteq{\mathrm{Gproj}}^{\leq k}(\Gamma)$ and so
@@
On the other hand, by [2, Lemmas 2.3 $(a)$ and 2.4 $(a)$], for every $Y\in T_{k}^{\perp}$, there is a short exact sequence
@@
where $T^{\prime}_{0},\cdots,T^{\prime}_{k-1}\in{\rm{add}}(T_{k})$. Clearly $K_{k-1}\in X^{\perp_{[k+1,\infty)}}$ and $T^{\prime}_{0},\cdots,T^{\prime}_{k-1}\in{\rm{add}}(T_{k})\subseteq X^{\perp}$. Hence, a dimension shifting argument shows that $Y\in X^{\perp}$. Therefore, $X\in{{}^{\perp}(T_{k}^{\perp})}$. But by [2, Lemmas 2.3 $(a)$ and 2.4 $(a)$] there is a short exact sequence
@@
Among the above mentioned cotorsion pairs, one of them has a very nice characteristic and its cotorsion class is a torsion-free class of a hereditary torsion pair. This fact relates higher Auslander-Gorenstein Algebras to the notion of torsion-cotorsion triples introduced by Bauer, Botnan, Oppermann and Steen [11] and gives us a nice description of them.
@@
A torsion cotorsion triple in an abelian category $\mathcal{A}$ is a triple of subcategories $(\mathcal{T},\mathcal{F},\mathcal{D})$ such that the pair $(\mathcal{T},\mathcal{F})$ is a torsion pair and the pair $(\mathcal{F},\mathcal{D})$ is a cotorsion pair, see [11, Page 29 before Theorem 2.33].
Theorem 5.3 (Dual of [11, Theorem 2.35]).
@@
This theorem is the dual version of [11, Theorem 2.35] and for the convenience of the reader we give the proof.
By [5, Proposition IV.1.9], we know that the inverse of Auslander-Reiten translation $\tau^{-}$ defines an equivalence
$\overline{{\rm{{mod}}}}(\Gamma)$$\underline{{\rm{{mod}}}}(\Gamma).$$\scriptstyle{\tau^{-}=\mathrm{Tr}\mathbf{D}}$
@@
where the middle equivalence follows from Auslander-Reiten formula because the injective dimension of every element of $\mathcal{D}$ by Bauer-Botnan-Oppermann-Steen Correspondence is at most one, see [42, Corollary III.6.4] and [11, Theorem 2.34]. ∎
@@
Since $\Gamma$ is an $n$-minimal Auslander-Gorenstein algebra, the triple $(^{\perp_{0}}\Gamma,{\mathrm{Gproj}}^{\leq n}(\Gamma),\mathrm{inj}^{\leq 1}(\Gamma))$ is a torsion cotorsion triple, see Corollary 3.7 and the arguments at the beginning of this section. Hence, by [11, Theorem 2.33], we have the first equivalence. That is
@@ 6. Higher Auslander Algebras
$\Gamma$ is also called a higher Auslander algebra, if there is a positive integer $n$ such that $\Gamma$ is an $n$-Auslander algebra. Iyama introduced these algebras as a generalization of classical Auslander algebras [28, Page 3].
@@
For $n\geq 2$, the conditions $\mathrm{proj}^{\leq n}(\Gamma)\cap\mathrm{inj}(\Gamma)\subseteq\mathrm{proj}(\Gamma)$ and ${{}^{\perp_{0}}\Gamma}\subseteq{{}^{\perp_{n}}\Gamma}$ in Theorem 6.1(b) are necessary. In fact, take an Artin algebra of global dimension n1n-1 which is not an $n$-Auslander algebra. Then $\mathrm{proj}^{\leq n-1}(\Gamma)={\rm{{mod}}}(\Gamma)$ is an abelian category, but by Theorem 6.1, the conditions that $\mathrm{proj}^{\leq n}(\Gamma)\cap\mathrm{inj}(\Gamma)\subseteq\mathrm{proj}(\Gamma)$ or ${{}^{\perp_{0}}\Gamma}\subseteq{{}^{\perp_{n}}\Gamma}$ don’t hold.
@@
$\begin{array}[]{rcl}\mathcal{M}&=&\{X\in{\rm{{mod}}}(\Lambda)\ \ |\ \ {\rm{Ext}}^{i}_{\Lambda}(X,\mathcal{M})=0,\ \forall\ 0<i<n\}\\ &=&\{X\in{\rm{{mod}}}(\Lambda)\ \ |\ \ {\rm{Ext}}^{i}_{\Lambda}(\mathcal{M},X)=0,\ \forall\ 0<i<n\}.\end{array}$
@@
The following example gives a better perspective about some results of the paper. In the following, for every vertex $i$ of the given quiver $Q$, the corresponding simple, projective, and injective object will be denoted by $S(i)$, $P(i)$, and $I(i)$, respectively. For more details about quiver representations, see for example [42].
@@
    By [42, Theorem I.10.5], the Nakayama algebra $\Gamma$ is of finite representation type with $2n+3$ indecomposable $\Gamma$-modules. The following table gives us more information about their homological dimensions.
@@
    $0\longrightarrow P(1)\longrightarrow P(n+1)\longrightarrow S(n+1)\longrightarrow 0.$
@@
Let $R$ be a commutative artinian ring and $n$ be a positive integer. In 2017, Iyama and Jasso extended higher Auslander correspondence from Artin $R$-algebras of finite representation type to dualizing $R$-varieties. By their result, i.e. higher Auslander correspondence for dualizing $R$-varieties, we know that an Artin $R$-algebra $\Gamma$ is an $n$-Auslander algebra if and only if the category of all finitely generated projective $\Gamma$-modules is an $n$-abelian category [29, Theorem 1.2]. But by some results due to Beligiannis [12], even for $n$-minimal Auslander-Gorenstein algebras, one can see that the category of all finitely generated Gorenstein projective $\Gamma$-modules is not an $n$-abelian category in general. Hence, based on our results, there is a difference between the abelianness of ${\mathrm{Gproj}}^{\leq n-1}(\Gamma)$ and the $n$-abelianness of ${\mathrm{Gproj}}(\Gamma)$ in general.
Acknowledgments: It is our great pleasure to thank many reseachers for their insightful suggestions and remarks, which include, but not restricted to, Zhaoyong Huang, Xiaojin Zhang, Bin Zhou, and Shijie Zhu.